# Bayesian Optimization of Laser-Wakefield Acceleration via Spectral Pulse Shaping

### v1

@@ V.1 Reference result: Gaussian pulse
The standard way to model LWFA in the modeling community is to use a Gaussian laser pulse, both transversely and longitudinally. We have already enhanced the fidelity of our result by modeling the transverse profile with a Jinc profile, but typically in most studies that do that a Gaussian longitudinal profile is still used despite not matching experiments. We will provide here a reference result to which we will compare our BO-enhanced results. In this case, we based our Gaussian mode on what is typically used to model the BELLA PW, that is 52 $\mu m$, $\tau=92\ fs$, and an energy of 30 J, which gives an approximate intensity of $I_{0}=4.53\times 10^{16}$ W/cm2 or normalized intensity of $a_{0}=1.46$, placing this design in an intermediate place between the quasi-linear and blowout regimes. Using a Gaussian pulse, we achieve $E_{m}=9.36$ GeV, $Q=10.31$ pC, and $\sigma_{E}=$ 2.96 % after 47.85 cm, which is within expectations per other simulation studies.benedetti2018
@@ V.2 Grid scan with respect to spectral shaping
To elucidate how spectral shaping affects the drive laser and baseline metrics of electron bunch mean energy and charge, we present grid scans of GDD, TOD and FOD in Fig. 3. In the first column, we plot the laser pulse shapes relative to spectral shaping and normalized them to the transform-limited pulse shape and compared them to an idealized pulse with a Gaussian longitudinal profile with a transverse Jinc profile. In the second column, we show the mean energy of the accelerated electron bunch and in the third column the amount of charge in the bunch. We can see that the Gaussian performs well overall, reaching the highest energy and then an intermediate level of charge. As we go higher in spectral content, even while increasing the bounds by an order of magnitude, the smaller the effect on the laser pulse and accelerated bunch properties. However, the complexity is further extended so that the spectral components can mix, in particular GDD and TOD, thus further expanding the parameter space. Mixing spectral components will allow us to exceed the performance of an idealized Gaussian pulse. To help us explore this large parameter space, we will use BO.
@@ V.3 Bayesian Optimization with respect to spectral shaping
The initial thrust of this study was to use spectral shaping, i.e., manipulating GDD, TOD, and FOD, to maximize the mean energy of the accelerated electron bunch, $E_{m}$. In this case and all subsequent cases, we only consider electrons with 0.1 GeV in energy or more. This study entailed several iterations of progressive complexity and breadth to find a result that achieved the target goals. Initially, we started with the use of BO to minimize a simple objective function of $f_{\text{obj}}=-E_{m}$. This proved to be insufficient because while $E_{m}$ was readily enhanced to $\gtrsim$ 20 GeV, it was done so at severe loss of charge, i.e., $\ll$ 1 pC levels.
To compensate for this, we implemented a composite objective function on which BO will operate of the following form:
@@
where again $Q$ is the charge in the electron bunch, $\sigma_{E}=\Delta E/E_{m}$ is the energy spread in the accelerated electrons and $\Delta E$ is defined as the standard deviation about the electron energy peak, and $a$, $b$, and $c$ are hyperparameters used to adjust the predictive performance of the GP so that it best reproduces the data prior to selecting new points to sample in parameter space. For this initial investigation, the hyperparameters were set to $a=0$, $C_{Q}=10^{14}$, and $C_{E}=C_{\sigma}=b_{Q}=b_{E}=b_{\sigma}=1$. These were manually tuned during the study. Similarly, this initial study only simulated the interaction up to 40 cm, while dephasing can occur as late as 60 cm in the following studies. The initial BO scan can be found in Fig.4, where we have plotted a) the simulation iteration # as the GP explores the parameter space, minimizing uncertainty and the objective function, and then b) GDD, c) TOD, and d) FOD parameters during optimization.
@@
Figure 5: The laser profiles and associated electron energy spectra for a,b) the Gaussian from BO scan #1 (orange) and #2 (blue), c,d) the peak performer for BO scan #1, e,f) the high energy case (HE), and g,h) the high charge (HQ) case. The laser profiles are normalized to the transform-limited intensity.
@@ V.4 Bayesian Optimization with respect to spectral shaping, laser, and plasma parameters
Although we were able to enhance the charge content of the accelerated bunch with BO, we did so at the expense of mean energy. To see if we can further optimize the system, we extend our GP parameter space from 3 to 8 dimensions, where we now include the central wavelength $\lambda_{c}[nm]$, injection distance $z_{\text{inj}}[cm]$, dopant layer thickness $\Delta z[cm]$, dopant fraction $f_{\text{dop}}[\%]$, and longitudinal density slope $L_{\text{slope}}[m^{-1}]$, where the base and boundary values are listed in Table 1.
@@
One of the disadvantages of GPs for regression or optimization is that they do not scale well with larger dimensions. For example, in Fig. 6, when we increase the parameter space from 3 to 8, the GP surrogate model is less capable of reproducing the training data without further calibration, thus the hyperparameters in Eq. 6. These were manually calibrated several times while generating and training new batches of simulations, where in the end we ran 193 simulations with coefficients $a=2$, $C_{Q}=10^{15}$, $C_{E}=C_{\sigma}=1$, and $b_{Q}=b_{E}=b_{\sigma}=5$. This was done in steps, initially starting on a gradient scan per each parameter, i.e., 8 simulations, then subsequently batches of individual simulations for the next 8 simulations varying all parameters, batches of three simulations for the next 42, and then 130 simulations in batches of 10. A unique advantage of GPs is their ability to select and train on new points in parallel, unlike gradient-based methods.
@@
The parameter space sampled with respect to the output metrics is presented in Fig. 7, where we show the $E_{m}$ vs. $Q$ slice. We can see that the parameter space sampled with just spectral components (orange) is much smaller than that sampled using an 8-dimensional sampling space (blue), which is to be expected. The baseline Gaussian pulse is also plotted (red). Depicted is an estimate of the Pareto front (dashed), where we plot $Q(E_{m}-5)=3\times 10^{3}$, which takes advantage of the simple intuition that the product $QE_{m}\approx const.$ is often approximately conserved in ideal circumstances tzoufras2008.
@@
Figure 6: Comparison of 3 parameter training on initial dataset versus 8 parameter training on a new dataset, where the sample points are circles and the GP surrogate prediction are crosses.
@@
Figure 7: Parameter space sampling with respect to output metrics $E_{m}$ and $Q$. Presented are the initial sampling with respect to just spectral parameters GDD, TOD, and FOD (orange), and then the larger sampling with respect to also $\lambda_{c}$, $z_{\text{inj}}$, $\Delta z$, $f_{\text{mix}}$, $L_{\text{slope}}$ (blue). Plotted also is an approximate Pareto front (dashed). Reference Gaussian is presented as well (star), for spectral-only, 3-dimensional BO case (orange), and for the 8-dimensional BO case (blue).
@@
A matrix plot of the Spearman rank correlation is presented in Fig. 8. The Spearman rank correlation, $\rho=r\left(\text{rank}(X),\text{rank}(Y)\right)$ and $r=\text{cov}(X,Y)/\sigma_{X}\sigma_{Y}$, measures monotonic associations both linear and nonlinear, making it less sensitive to outliers and invariant to monotonic transformations, e.g., $\log_{10}(Q)$. The strongest correlation is between $f_{\text{dop}}$ and $E_{m}$, where the energy drops as we increase the mix fraction. This correlates to beamloading, wherein a higher electron bunch charge reduces the amplitude of the accelerating field, and thereby achieves lower energies. Based on this figure and additional analysis, it appears that FOD, $\lambda_{c}$, and injection distance $z_{\text{inj}}$ have little effect on $Q$ or $E_{m}$. Higher levels of dopant fraction, GDD, and $L_{\text{slope}}$ result in case examples with the highest values of $Q$ but lowest values of $E_{m}$.
@@
Figure 8: $Q$ versus $E_{m}$ parameter space with respect to the input parameters. A correlation matrix of inputs versus outputs with respect to Spearman rank correlation.
@@ VI High-performing, BO-enabled case examples
With only spectral shaping we could significantly increase the charge relative to the Gaussian case but not the mean energy as much without losing charge in turn. By incorporating the laser wavelength, injection range, and channel plasma conditions into the BO loop, we found new optima for both $E_{m}$ and $Q$ without severely sacrificing one or the other. The optimization path was somewhat skewed by the objective function, thus the clustering of points near $Q\sim 200$ pC and $E_{m}\sim$ 7 GeV. We will investigate several high performers and how they differ from the baseline Gaussian case.
Based on the eight-dimensional objective function and its associated coefficients, the peak performer selected, which we designate the High Energy case (HE), has the following setup conditions found in Table 2. These can be compared to the baseline parameters that informed the 3-dimensional optimization scan and the reference Gaussian pulse case. Several of the HE parameters approach the boundaries of the initial sampling range, notably $\lambda_{c}$, FOD, and $z_{\text{inj}}$, which may suggest pushing those limits further. The HE metrics are $E_{m}=15.3$ GeV, $Q=85.4$ pC, and $\sigma_{E}=7.7$ %. The associated laser pulse and electron energy spectrum at peak $E_{m}$ can be found in Fig. 5(c). Another notable case is the fourth highest performing simulation, the High Charge case (HQ), whose input parameters can be found in Table. 3 and the associated laser and energy spectrum in Fig. 5(d). The output metrics for this simulation are $E_{m}=8.43$ GeV, $Q=597.3$ pC, and $\sigma_{E}=6.1$ %.
@@
There are several physical reasons why these pulse shapes perform better than the Gaussian pulse. One is that the spectral content of the laser reduces beating in the laser intensity as it propagates through the plasma, resulting in a more stable accelerating electric field. This can be seen in Fig. 9.a). What is especially noticeable is that shaped pulses do not oscillate significantly relative to the Gaussian laser pulse.benedetti2012; benedetti2015 Likewise, there is an appreciable increase in the laser intensity over time, particularly for the HQ case. This means that $E_{z}$ will increase accordingly, further enhancing the acceleration of the electron bunch and likely compensating for any beamload due to the very high bunch charge ($\gtrsim$ 500 pC). This is illustrated by plots of $E_{m}$ and $Q$ with respect to time as shown in Fig. 9.b), with significant charge injection (several hundred pC), in all three cases. Unlike the HE and HQ cases, the Gaussian loses much of the charge during the initial beating of the laser mode. If we focus earlier in time in Fig. 9.c) we observe that not only is the Gaussian pulse losing charge over time, with a precipitous drop around 4.5 cm, but the HQ is actually increasing in the amount of charge injected despite periodic losses due to beating. The HE case injects less charge overall than the Gaussian or HQ case but maintains it throughout until depletion occurs. We can further visualize the electron bunch relative to the bubble within which it is being accelerated, as seen in Fig. 10. Here, the bubble is defined simply as the first bucket in the wake where the electric field is negative. This illustrates how the bunch leaves and reenters the bubble during oscillations correlates with the beating in laser intensity $a_{0}$.
Figure 9: a) Comparison of peak $a_{0}$ as a function of propagation distance for the baseline Gaussian (red), HE case (orange), and HQ case (red) and then b) $E_{m}$ and $Q$ as a function of time. In sub-figure c) we focus in on $E_{m}$ and $Q$ earlier in time, where the dopant region cutoff is illustrated (vertical).
@@ VII Conclusions
In this work, we used Bayesian optimization to investigate the effects of spectral shaping and plasma channel parameters on key LWFA figures of merit, i.e., the mean energy of the accelerated electron bunch $E_{m}$, charge $Q$ and energy spread $\sigma_{E}$. Spectral shaping alone can have an effect on LWFA, particularly GDD and TOD, but not enough to supersede a baseline Gaussian reference simulation. Incorporating additional plasma channel parameters, such as dopant length, dopant fraction, and a linear longitudinal density slope, allowed for significant improvement. The best performer achieved 15 GeV energies within the current parameter boundaries without sacrificing high charge. In fact, not only can we increase energy with a noticeable increase of charge (from 10 to  500 pC), but we can also achieve comparable energies but with near nC charge content levels.
@@
This is a relatively ad hoc application of BO to LWFA and can be expanded further. A more realistic laser modeling approach could be incorporated, i.e., modeling the effects of amplification, compression, focusing, deformable mirrors, etc. Similarly, even more realistic plasma channels could be modeled. For example, a linear density slope was applied across the entire plasma channel in this study, while a realistic, ionized gas jet would have different channel transverse curvatures and profiles based on the local density longitudinally. In addition, multi-fidelity BO can be used as well as more sophisticated approaches to multi-objective BO.

### v2

@@ V.1 Reference result: Gaussian pulse
The standard way to model LWFA in the modeling community is to use a Gaussian laser pulse, both transversely and longitudinally. We have already enhanced the fidelity of our result by modeling the transverse profile with a Jinc profile, but typically in most studies that do that a Gaussian longitudinal profile is still used despite not matching experiments. We will provide here a reference result to which we will compare our BO-enhanced results. In this case, we based our Gaussian mode on what is typically used to model the BELLA PW, that is 52 $\mu m$, $\tau=92\ fs$, and an energy of 30 J, which gives an approximate intensity of $I_{0}=4.53\times 10^{16}$ W/cm2 or normalized intensity of $a_{0}=1.46$, placing this design in an intermediate place between the quasi-linear and blowout regimes. Using a Gaussian pulse, we achieve $E_{m}=9.36$ GeV, $Q=10.31$ pC, and $\sigma_{E}=$ 2.96 % after 47.85 cm. WarpX simulations using these parameters were calibrated to and closely matched preliminary INF&RNO simulations.benedetti2018
@@ V.2 Grid scan with respect to spectral shaping
To elucidate how spectral shaping affects the drive laser and baseline metrics of electron bunch mean energy and charge, we present grid scans of GDD, TOD and FOD in Fig. 3. In the first column, we plot the laser pulse shapes relative to spectral shaping and normalized them to the transform-limited pulse shape and compared them to an idealized pulse with a Gaussian longitudinal profile with a transverse Jinc profile. In the second column, we show the mean energy of the accelerated electron bunch and in the third column the amount of charge in the bunch. We can see that the Gaussian performs well overall, reaching the highest energy and then an intermediate level of charge. As we go higher in spectral content, even while increasing the bounds by an order of magnitude, we have a smaller effect on the laser pulse and accelerated bunch properties. However, the complexity is further extended so that the spectral components can mix, in particular GDD and TOD, thus further expanding the parameter space. Modifying multiple spectral components at once will allow us to exceed the performance of an idealized Gaussian pulse. To help us explore this large parameter space, we will use BO.
@@ V.3 Bayesian Optimization with respect to spectral shaping
The initial thrust of this study was to use spectral shaping, i.e., manipulating GDD, TOD, and FOD, to maximize the mean energy of the accelerated electron bunch, $E_{m}$. In this case and all subsequent cases, we only consider electrons with 0.1 GeV in energy or more. This study entailed several iterations of progressive complexity and breadth to find a result that achieved the target goals. Initially, we started by using BO to minimize a simple objective function of $f_{\text{obj}}=-E_{m}$. This proved to be insufficient because while $E_{m}$ was readily enhanced to $\gtrsim$ 20 GeV, it was done so at severe loss of charge, i.e., $\ll$ 1 pC levels.
To compensate for this, we implemented a composite objective function on which BO operated of the following form:
@@
where again $Q$ is the charge in the electron bunch, $\sigma_{E}=\Delta E/E_{m}$ is the energy spread in the accelerated electrons and $\Delta E$ is defined as the standard deviation about the electron energy peak, and $a$, $b$, and $C$ are hyperparameters used to adjust the predictive performance of the GP so that it best reproduces the data prior to selecting new points to sample in parameter space. For this initial investigation, the hyperparameters were set to $a=0$, $C_{Q}=10^{14}$, and $C_{E}=C_{\sigma}=b_{Q}=b_{E}=b_{\sigma}=1$. These were manually tuned during the study. Similarly, this initial study we only simulated the interaction up to 40 cm, while dephasing can occur as late as 60 cm in the following studies. The initial BO scan can be found in Fig.4, where we have plotted a) the simulation iteration # as the GP explores the parameter space, minimizing uncertainty and the objective function, and then b) GDD, c) TOD, and d) FOD parameters during optimization.
@@
Figure 5: The laser profiles and associated electron energy spectra for a,b) the Gaussian with respect to the simulation run distance for BO scan #1 (orange, 40 cm) and #2 (blue, 75 cm), c,d) the peak performer for BO scan #1, e,f) the high energy (HE) and g,h) the high charge (HQ) cases from BO scan #2. The laser profiles are normalized to the transform-limited intensity.
@@ V.4 Bayesian Optimization with respect to spectral shaping, laser, and plasma parameters
Although we were able to enhance the charge content of the accelerated bunch with BO, we did so at the expense of mean energy. To see if we can further optimize the system, we extend our GP parameter space from 3 to 8 dimensions as well as extend the range of TOD and FOD per Table 1, where we now include the central wavelength $\lambda_{c}[nm]$, injection distance $z_{\text{inj}}[cm]$, dopant layer thickness $\Delta z[cm]$, dopant fraction $f_{\text{dop}}[\%]$, and longitudinal density slope $L_{\text{slope}}[m^{-1}]$.
@@
One of the disadvantages of GPs for regression or optimization is that they do not scale well with larger dimensions. For example, in Fig. 6, when we increase the parameter space from 3 to 8 but use the same hyperparameters per Sec. V.3, the GP surrogate model is less capable of reproducing the training data without further calibration, thus the inclusions of hyperparameters in Eq. 6. These were manually calibrated several times while generating and training new batches of simulations, where in the end we ran 193 simulations with coefficients $a=2$, $C_{Q}=10^{15}$, $C_{E}=C_{\sigma}=1$, and $b_{Q}=b_{E}=b_{\sigma}=5$. This was done in steps, initially starting with a gradient scan with respect to each parameter, i.e., 8 simulations, then subsequently individual simulations for the next 8 simulations, varying all parameters, batches of three simulations for the next 42, and then 130 simulations in batches of 10. A unique advantage of GPs is their ability to select and train on new points in parallel, unlike gradient-based methods.
@@
The parameter space sampled with respect to the output metrics is presented in Fig. 7, where we show the $E_{m}$ vs. $Q$ slice. We can see that the parameter space sampled with just spectral components (orange) is much smaller than that sampled using an 8-dimensional sampling space (blue), which is to be expected. The baseline Gaussian pulse is also plotted for both datasets (stars), where the simulation range was increased from 40 cm to 75 cm. Depicted is an estimate of the Pareto front (dashed), where we plot $Q(E_{m}-5)=3\times 10^{3}$, which takes advantage of the simple intuition that the product $QE_{m}\approx const.$ is often approximately conserved in ideal circumstances tzoufras2008.
@@
Figure 6: Comparison of 3 parameter training on initial dataset versus 8 parameter training on a new dataset, including the sample points (circles) and the GP surrogate prediction (crosses).
@@
Figure 7: Parameter space sampling with respect to output metrics $E_{m}$ and $Q$. Presented are the initial sampling with respect to just spectral parameters GDD, TOD, and FOD (orange), and then the larger sampling with respect to also $\lambda_{c}$, $z_{\text{inj}}$, $\Delta z$, $f_{\text{mix}}$, $L_{\text{slope}}$ (blue). Plotted also is an approximate Pareto front (dashed). Reference Gaussian is presented as well, for the spectral-only, 3-dimensional BO case (orange star), and for the 8-dimensional BO case (blue star).
@@
A matrix plot of the Spearman rank correlation is presented in Fig. 8. The Spearman rank correlation, $\rho=r\left(\text{rank}(X),\text{rank}(Y)\right)$ with $r=\text{cov}(X,Y)/\sigma_{X}\sigma_{Y}$, measures monotonic associations both linear and nonlinear, making it less sensitive to outliers and invariant to monotonic transformations, e.g., $\log_{10}(Q)$. The strongest correlation is between $f_{\text{dop}}$ and $E_{m}$, where the energy drops as we increase the mix fraction. This correlates to beamloading, wherein a higher electron bunch charge reduces the amplitude of the accelerating field, and thereby achieves lower energies. Based on this figure and additional analysis, it appears that FOD, $\lambda_{c}$, and injection distance $z_{\text{inj}}$ have little effect on $Q$ or $E_{m}$. Higher levels of dopant fraction, GDD, and $L_{\text{slope}}$ result in case examples with the highest values of $Q$ but lowest values of $E_{m}$.
@@
Figure 8: $Q$ and $E_{m}$ variance with respect to the input parameters. A correlation matrix of inputs versus outputs with respect to Spearman rank correlation.
@@ VI High-performing, BO-enabled case examples
With only spectral shaping we could significantly increase the charge relative to the Gaussian case but not the mean energy significantly without losing charge in turn. By incorporating the laser wavelength, injection range, and channel plasma conditions into the BO loop, we found new optima for both $E_{m}$ and $Q$ without severely sacrificing one or the other. The optimization path was somewhat skewed by the objective function, thus the clustering of points near $Q\sim 200$ pC and $E_{m}\sim$ 7 GeV. We will investigate several high performers and how they differ from the baseline Gaussian case.
Based on the eight-dimensional objective function and its associated coefficients, the peak performer selected, which we designate the High Energy case (HE), has the following setup conditions found in Table 2. These can be compared to the baseline parameters that informed the 3-dimensional optimization scan and the reference Gaussian pulse case. Several of the HE parameters approach the boundaries of the initial sampling range, notably $\lambda_{c}$, FOD, and $z_{\text{inj}}$, which may suggest pushing those limits further. The HE metrics are $E_{m}=15.3$ GeV, $Q=85.4$ pC, and $\sigma_{E}=7.7$ %. The associated laser pulse and electron energy spectrum at peak $E_{m}$ can be found in Fig. 5(e,f). Another notable case is the fourth highest performing simulation, the High Charge case (HQ), whose input parameters can be found in Table. 3 and the associated laser and energy spectrum in Fig. 5(g,h). The output metrics for this simulation are $E_{m}=8.43$ GeV, $Q=597.3$ pC, and $\sigma_{E}=6.1$ %.
@@
There are several physical reasons why these pulse shapes perform better than the Gaussian pulse. One is that the spectral content of the laser reduces beating in the laser intensity as it propagates through the plasma,leemans2014; djordjevic2018 resulting in a more stable accelerating electric field. This can be seen in Fig. 9.a). It is noticeable that spectrally shaped pulses do not oscillate significantly relative to the Gaussian laser pulse.benedetti2012; benedetti2015 Likewise, there is an appreciable increase in the laser intensity over time, particularly for the HQ case. This means that $E_{z}$ will increase accordingly, further enhancing the acceleration of the electron bunch and likely compensating for any beamloading due to the high bunch charge ($\gtrsim$ 500 pC). This is illustrated by plots of $E_{m}$ and $Q$ with respect to time as shown in Fig. 9.b), with significant charge injection (several hundred pC), in all three cases. Unlike the HE and HQ cases, the Gaussian loses much of the charge during the initial beating of the laser mode. If we focus earlier in time in Fig. 9.c), we observe that not only is the Gaussian pulse losing charge over time, with a precipitous drop around 4.5 cm, but the HQ is increasing in the amount of charge injected despite periodic losses due to beating given the extended dopant region (vertical line). The HE case injects less charge overall than the Gaussian or HQ case but maintains it throughout until depletion occurs. We can further visualize the electron bunch relative to the bubble within which it is being accelerated, as seen in Fig. 10. Here, the bubble is defined simply as the first bucket in the wake where the electric field is negative. This illustrates how the bunch leaves and reenters the bubble during oscillations correlates with the beating in laser intensity $a_{0}$.
Figure 9: a) Comparison of peak $a_{0}$ as a function of propagation distance for the baseline Gaussian (red), HE case (orange), and HQ case (red) and then b) $E_{m}$ and $Q$ as a function of time. In sub-figure c) we focus in on $E_{m}$ and $Q$ earlier in time, where the dopant region cutoff is illustrated (vertical lines).
@@ VII Conclusions
In this work, we used Bayesian optimization to investigate the effects of spectral shaping and plasma channel parameters on key LWFA figures of merit, i.e., the mean energy of the accelerated electron bunch $E_{m}$, charge $Q$ and energy spread $\sigma_{E}$. Spectral shaping alone can have an effect on LWFA, particularly GDD and TOD, greatly increasing $Q$ but not appreciably increasing the bunch energy $E_{m}$. Incorporating additional laser and plasma channel parameters, such as injection distance, dopant length, dopant fraction, and a linear longitudinal density slope, allowed for significant improvement. The best performer achieved 15 GeV energies within the current parameter boundaries without sacrificing high charge. In fact, not only can we increase energy with a noticeable increase of charge (from 10 to  500 pC), but we can also achieve comparable energies but with approaching nC charge content levels while keeping laser energy constant.
@@
This is a relatively ad hoc application of BO to LWFA and can be expanded further. A more realistic laser modeling approach could be incorporated, i.e., modeling the effects of amplification, compression, focusing, deformable mirrors, etc. Similarly, even more realistic plasma channels could be modeled. For example, a linear density slope was applied across the entire plasma channel in this study, while a realistic, ionized gas jet would have different channel transverse curvatures and profiles based on the local density longitudinally. In addition, multi-fidelity BO can be used,gratiet2013 and recent work has demonstrated fully multi-objective and multi-fidelity BO for LWFA using hypervolume-based and trust-aware acquisition policies that reduce the cost of exploring Pareto-optimal beam settings,irshad2023 an approach that could be extended to spectral shaping.